OpenStudy (anonymous):

Use graphical methods to solve the linear programming problem. 2x + 6y ≥ 30. 4x + 2y ≥ 20. x ≥ 0. y ≥ 0. Minimize P = 0.18x + 0.12y A. Minimum of 1.86 when x = 9 and y = 2 B. Minimum of 1.08 when x = 4 and y = 3 C. Minimum of 1.02 when x = 3 and y = 4 D. Minimum of 1.2 when x = 4 and y = 4

OpenStudy (tkhunny):

You should graph on the same coordinate axes: y = 5 - x/3 y = 10 - 2x This will give you a good start. (Where did those come from?)